[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Sep 21, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Sep 21 12:02:32 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. The AWIPS LP1 printer is complaining about waste toner box and thus  
nothing will print. This issue has been occurring all year. An NCO Help  
Desk ticket has been submitted by TSB.

2. Dvorak classification/fix images would not print for Karl, TD13 or Paine  
from either East Pacific or Atlantic desks, even after trying several  
times. Had to manually print each image. TSB will investigate.

3. Altimeter data did not show up in N-AWIPS all night. [Does this come in  
locally through Mike Brennan's box?] Also had NMAP WW3 procedure bombed on  
me at beginning of shift and thought both issues might be associated with  
incoming data problems to Mike B's box. However, these WW3 procedures now  
load since 09Z for 00Z runs. TSB will investigate.

Most of this processing has been migrated from Mike Brennan's  
workstation 'pinfish' to NHC's comp01/03 servers, for example:

RFC #181 – get_ascatb_nhc v2.0 - Migrate script that wget's ASCAT-B data  
from the NESDIS 'manati' system to redundant NHC compute servers. This  
transition will facilitate troubleshooting and make the ingest of ASCAT-B  
data more robust/less prone to outages. Implemented on Monday, June 27,  
2016 at 1500Z (11 am EDT).
HSU:
1. ATCF issues: The Pacific desk ATCF workstation was having some screen  
refresh issues when windows were being rendered. Senior Hurricane  
Specialist Stacy Stewart rebooted the workstation, which corrected the  
problem.
Other items:
1. The limit for the number of processes allowed to user 'naprod' on NHC's  
primary compute server (nhc-lx-comp01) has been reset to the default value  
of 1024 by NCO Linux SysAdmin Curt Steinmetz.

2. IBM will apply a patch to fix the shared library/undefined symbols  
problem on the WCOSS supercomputing clusters, then NCO will take Gyre and  
Surge for a parallel production test starting with the 12z cycle on  
Thursday, September 22. TSB Developer Monica Bozeman is fixing  
NHC's 'prod2dev' rsync scripts to make sure the HWRF/GFDL guidance data  
reaches both the production and development machines.

3. Missing Recon Data in NMAP [from the SDM's log]:

1818Z to 1827Z - NHC's Monica Bozeman reported that recon aircraft
data is unavailable in NMAP...per Monica, NHC's Robbie is
working with WPC in an exercise to backup NHC operations, in
which recon data is needed to effectively complete the exercise.
Notified Dataflow about this.

2002Z - E-mail from NHC's Michael Brennan: "Yes, when I was
the SOO at HPC back in 2007-2008 I implemented the script that
created GEMPAK format ship files from the recon HDOBS. I think
it used to run on my workstation locally there. I'm not sure
what became of it after I left, but perhaps someone in DTB would
know". (RJS)

2011Z - Alan Robson will investigate the issue tomorrow (Sept. 21). (RJS)

4. WINDSAT data outage [from the SDM's log]:

1118Z - Big Brother is showing persistent yellow flags on the
vm-lnx-ncosrv mirror for missing Windsat data. Emailed Paul and
Qi about this outage.

0150Z - Flags continue (RJS)
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160921/a4e6f378/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list